I purchase 18"x18" tiles and they are sold 6 to a carton, how many cartons are needed for a 12'x9' floor?
Answer by checkley77(12844)

12*9=108 ft^2 is the floor area.
18*18=1.5*1.5=2.25 ft^2 for each tile
6*2.25=13.5 ft^2 per carton
108/13.5=8 cartons of tile will be needed.